OpenStudy (anonymous):

Evaluate the following expression using the values given: Find -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3. I need help !

OpenStudy (anonymous):

sole above expression you wii get -a2-b3 put the given values you will get the answ.

OpenStudy (whpalmer4):

\[-a^2-3b^3+c^2+2b^3-c^2\]Collecting like terms gives you \[-a^2-3b^3+2b^3+c^2-c^2 = -a^2-b^3\]Now put in your known values: \[-(3)^2-(2)^3 = \]Note well that the first term is the same as -1*3*3, NOT (-3)*(-3)!